OpenStudy (anonymous):

If a baseball player hits a baseball from 4 feet off the ground with an initial velocity of 64 feet per second, how long will it take the baseball to hit the ground? Use the equation h = -16t2 + 64t + 4.

OpenStudy (welshfella):

when the ball hits the ground h will = 0 so form an equation and solve for t

OpenStudy (anonymous):

So far I have this 64 +/- sqrt 3840 over 32. But I don't know what to do after that

OpenStudy (welshfella):

make it easier by taking out -4 -4(4t^2 - 16t - 1) = 0 now apply quadratic formula to the expression in the parentheses t =[ 16 +- sqrt(256 - 4*4 * -1) ] / 8

OpenStudy (welshfella):

t = (16 + 16.49) / 8 (neglect the negative root)

OpenStudy (anonymous):

im confused

OpenStudy (welshfella):

haven't you used the quadratic formula before?

OpenStudy (anonymous):

yes but I didn't get what you got

OpenStudy (welshfella):

i applied it to 4x^2 - 16x - 1 = 0

OpenStudy (welshfella):

- smaller numbers

OpenStudy (anonymous):

oh ok

OpenStudy (welshfella):

i divide through by -4 to make it easier

OpenStudy (welshfella):

this equation will have the same roots as the bigger one

OpenStudy (anonymous):

so the answer I got was 16+/- sqrt 17 over 2 is that what you got?

OpenStudy (welshfella):

(16 + sqrt272 ) / 8 = 4.06 secs

OpenStudy (anonymous):

my answer has to be in the quadratic formula

OpenStudy (welshfella):

oh ok so that's (16 + sqrt(16*17) ) / 8 = 16 + 4*sqrt17 ----------- 8 = 2 + 0.5 sqrt17

OpenStudy (anonymous):

ok thank you

OpenStudy (welshfella):

yw